OpenStudy (anonymous):

How to get from 5(2^(n-1) + 4 * 3^(n-1)) - 6(2^(n-2) + 4 * 3^(n-2)) to (5 * 2 - 6)2^(n-2) + (5 * 4 * 3 - 6 * 4)3^(n-2) by simplifying the initial expression?

OpenStudy (amistre64):

what have you worked out so far?

OpenStudy (anonymous):

Nothing this isn't the problem it's part of a solution and I'm years out of practice in math and I just have absolutely no idea how to get from that one piece to the next.

OpenStudy (amistre64):

factoring is key ... but so is some basic exponent rules

OpenStudy (amistre64):

distribute, collect like terms 2^this and 3^that and factor out what needs to be factored to turn it

OpenStudy (anonymous):

yeah i know that but i don't know HOW. So what, the first half of the first expression would be 10^(n-1) + 20 * 15^(n-1)? Is that correct? Then deal with the negative exponents or what

OpenStudy (amistre64):

as long as your willing to participate, im willing to help out ... your first reply was sufficient

OpenStudy (amistre64):

lets keep the actual multiplication in parts for the moment, jsut show the distribution itself 5(2^(n-1)) + 5(4)(3^(n-1)) - 6(2^(n-2)) + 6(4)(3^(n-2))) now we move things about since our end reult wants to be in terms of 2^ and 3^

OpenStudy (amistre64):

gathering like terms 5(2^(n-1))- 6(2^(n-2)) + 6(4)(3^(n-2)))+ 5(4)(3^(n-1)) now we are in position to factor out what needs to be facotred right?

OpenStudy (anonymous):

ok yeah

OpenStudy (anonymous):

no wouldn't it be - 6(2^(n-2)) - 6(4)(3^(n-2))) from your prev post

OpenStudy (anonymous):

since it's -6 * (4)(3^(n-2)))

OpenStudy (amistre64):

we know we want to factor out 2^(n-2) and 3^(n-2) [5(2^(n-1))/(2^(n-2)) - 6] 2^(n-2) + [6(4)+ 5(4)(3^(n-1))/3^(n-2) ] 3^(n-2)

OpenStudy (amistre64):

the like terms we are looking for is not 6, but looking at what we want ... its going to have to be the 3^ parts

OpenStudy (anonymous):

talking about the distribution 5(2^(n-1)) + 5(4)(3^(n-1)) - 6(2^(n-2)) + 6(4)(3^(n-2))) should be 5(2^(n-1)) + 5(4)(3^(n-1)) - 6(2^(n-2)) - 6(4)(3^(n-2))) instead or am i wrong

OpenStudy (amistre64):

lets look at this in the coding ..... \[5~[2^{n-1} + 4~3^{n-1}] - 6~[2^{n-2} + 4~3^{n-2}]\] distribute, and yes, it looks like i dropped that negative, my err \[5~2^{n-1} + 5~4~3^{n-1} - 6~2^{n-2} - 6~ 4~3^{n-2}\] reaarange the like terms \[5~2^{n-1}- 6~2^{n-2} + 5~4~3^{n-1} - 6~ 4~3^{n-2}\] factor out the required parts \[[5~\frac{2^{n-1}}{2^{n-2}}- 6]~2^{n-2} + [5~4~\frac{3^{n-1}}{3^{n-2}} - 6~ 4]~3^{n-2}\]

OpenStudy (amistre64):

exponent rule: b^n/b^m = b^(n-m) so that should get it to what you were looking for

OpenStudy (amistre64):

feels like the site is starting to breakdown again. notifs and posts are getting slowed down

OpenStudy (anonymous):

sok slow seems to be a better speed for my brain

OpenStudy (anonymous):

because how did you get from rearranging the like terms to "factor out the required parts" what happened there

OpenStudy (anonymous):

with the exponential terms parts

OpenStudy (amistre64):

the rearrange is just to get all the parts nest to each other that have the required factor that will be pulled out commutative property at work is all. the end results shows us what to work towards: () 2^ and () 3^ so i simply arranged them and pulled out the needed factor are you asking why a divided as i did in the last step?

OpenStudy (anonymous):

yes this whole thing factor out the required parts [5 2n−12n−2−6] 2n−2+[5 4 3n−13n−2−6 4] 3n−2

OpenStudy (anonymous):

well i just copy pasted it didn't paste right but the last part of your post

OpenStudy (amistre64):

lets make sure we are in step for the rearranging .... \[5~2^{n-1}- 6~2^{n-2} + 5~4~3^{n-1} - 6~ 4~3^{n-2}\] its good for you up to here right?

OpenStudy (amistre64):

refresh gets rid of the odd marks but then i was stuck in a death spiral refresh page load lol

OpenStudy (amistre64):

the concept i want to use here is that 1, times anything, changes nothing. 1(8) = 8 1(2) = 2 1(15) = 15 so, im thinking we need to multiply by a useful 'form' of 1 anything divided by itself (except 0/0) is equal to 1 sooo \[5~2^{n-1}- 6~2^{n-2} + 5~4~3^{n-1} - 6~ 4~3^{n-2}\] \[5~2^{n-1}~1- 6~2^{n-2} + 5~4~3^{n-1}~1 - 6~ 4~3^{n-2}\] \[5~2^{n-1}~\frac{2^{n-2}}{2^{n-2}}- 6~2^{n-2} + 5~4~3^{n-1}~\frac{3^{n-2}}{3^{n-2}} - 6~ 4~3^{n-2}\] and factor out the required parts \[[5~\frac{2^{n-1}}{2^{n-2}}- 6]~2^{n-2}~+~[5~4~\frac{3^{n-1}}{3^{n-2}} - 6~ 4]~3^{n-2}\]

OpenStudy (amistre64):

the last step is just simplifying the fractions .... using the exponent rule \[[5~\frac{2^{n-1}}{2^{n-2}}- 6]~2^{n-2}~+~[5~4~\frac{3^{n-1}}{3^{n-2}} - 6~ 4]~3^{n-2}\] \[[5~2^{n-1-n+2}- 6]~2^{n-2}~+~[5~4~3^{n-1-n+2} - 6~ 4]~3^{n-2}\] \[[5~2^{1}- 6]~2^{n-2}~+~[5~4~3^{1} - 6~ 4]~3^{n-2}\] \[[5~2- 6]~2^{n-2}~+~[5~4~3 - 6~ 4]~3^{n-2}\]